What quantity of 60 per cent acid solution must be mixed with a 30 per cent solution to produce 540 mL of a 50 per cent solution
Answer by modi.kk91(20)

let at 60% ml=x
and at 30%ml=540-x

60%*x+30%*(540-x)=540*50
60x+16200-3ox=27000

x=360 at 60%
and 540-360=180 at30%